OpenStudy (anonymous):

write the equation of a line parallel to the given line but passing through the given point. y=2x-1; (2, -2)

OpenStudy (anonymous):

I think you messed up on the equation. Where's the x?

OpenStudy (anonymous):

Alright. y=2x-1 Plug in the numbers. -2=2(2)-b (Because you're trying to find a new y intersection to have a parallel line with the equation, the slope stays the same.) -2=4-b -6=-b 6=b y=2x-6